2) How many Liters each of 15% acid and 33% acid should be mixed to get 120 Liters of 21% acid?
Answer by ewatrrr(24785)

 
Hi,
mixed to get 120 Liters
Let x and (120-x) represent the amount of the 33% and 15% solutions respectively
Question states***
mixing to get 120 Liters of 21% acid?
.33x + .15(120-x) = .21*120L
Solving for x
.18x = .06*120L = 7.2L
x = 40L, the amount of the 33% Solution. 80L of the 15% solution
CHECKING our Answer***
.33*40 + .15*80 = 13.2 + 12 = 25.2 = .21*120